A heat engine operates with reservoirs at 650° C and 15° C, but it operates at 35% of its maximum possible efficiency.

A. What is the maximum efficiency?

B. What is the actual total efficiency?

C. If 1500 J of work is done by the heat engine, how much energy is extracted from the hot reservoir?

D. How much energy is exhausted?
Answer #1

a)

max efficiency

n = 1 - T2/T1 = 1 - (273 + 15)/ (273 + 650)

n = 0.688 = 68.8 %

b)

actual efficiency

n' = 0.688* 0.35 = 24.08 %

======

c)

n' = W/Q1

0.2408 = 1500 / q1

q1 = 6229.24 J

====

d)

exhausted = 6229.24 - 1500 = 4729.2 J

=====

Comment before rate in case any doubt, will reply for sure..